# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

### v2

@@ Abstract
In the fast-paced field of quantum computing, identifying the architectural characteristics that will enable quantum processors to achieve high performance across a diverse range of quantum algorithms continues to pose a significant challenge. Given the extensive and costly nature of experimentally testing different designs, this paper introduces the first Design Space Exploration (DSE) for quantum-dot spin-qubit architectures. Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal. ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute force approach while maintaining the same result quality. After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific and universal architectural features that provide optimal performance across the examined circuits. Our work demonstrates that the synergy between DSE methodologies and optimization algorithms can effectively be deployed to provide useful suggestions to quantum processor designers.
@@ I Introduction
As the field of quantum computing rapidly advances, different qubit technologies exhibit unique hardware and performance characteristics. It still remains to be seen which one (e.g., superconducting, trapped ions, quantum dots, photonics, defect-based on nitrogen-vacancy diamond centers) will succeed in scaling up quantum computing systems with high-quality qubits [59, 12]. Among them, spin qubits in quantum dots emerge as a compelling avenue for achieving scalability for practical quantum computation [12]. To this day, spin qubits are at an early stage in their development, with Intel’s Tunnel Falls chip currently boasting the highest count of twelve spin qubits [34]. Despite this, their inherent scalability advantages suggest that a robust two-dimensional design [68, 38, 8] could be scaled up relatively easily once fabrication techniques and quality improve up to a certain level. Designing a chip, however, involves multiple architectural design choices whose impact in the future can only be fully assessed through experimental studies post-scaling the technology. Exploring all these design possibilities simultaneously can be expensive, time-intensive, and impractical. Therefore, simulating how different design choices can affect performance in a range of quantum applications is crucial to assist the development. This process facilitates highlighting architectural insights that can guide the technology forward and will allow quantum researchers to make more informed decisions, streamline efforts, and hasten the development of these devices.
@@
Standard DSE processes [54, 43] will be impractical time-wise, especially for such large spaces, as they rely on a brute force approach to traverse and subsequently analyze the design space. Our approach, however, employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods designed to automate the DSE process of SpinQ, thus taking less time than the brute force approach with the same quality of results. The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum benchmark circuit.
Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing. Then, equipped with these insights, we move on to conduct a DSE analysis of $29,312$ spin qubit architectures and provide valuable insights into best design practices. Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits. Secondly, we find that the communication method with shuttle operations is better in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based methods rotations are preferred over shuttle-based. Finally, our universal architecture for all used quantum circuits suggests that the likelihood of success increases when prioritizing the parallelization of single-qubit gates instead of two-qubit gates. With these, we showcase that combining DSE methodologies with optimization algorithms effectively assists in forming beneficial recommendations for quantum processor designers for current and future technological stages.
The main contributions of this paper:
@@
    The new SpinQ, the first compiler and DSE framework for spin qubit architectures. In this version, we have updated the ESP formula to include operational, crosstalk, and decoherence-induced errors.
@@
The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the performance of the resulting architecture. Then, in Section III, we motivate the need for an automated Design Space Exploration analysis for spin qubit devices and formulate four research questions for this work. After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new DSE functionality of SpinQ. Another upgrade in SpinQ is the enhanced Estimated Success Probability metric, described in Section V, which is used as the Figure of Merit for our exploration. In Section VI, we introduce ArtA (Artificial Architect) and two new metrics to evaluate its performance. In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration. Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights for a universal architecture. We conclude our work and provide ideas for future directions in Section IX.
@@ II Challenges for scaling up spin qubit devices
Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing. The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high temperatures [76, 11, 32, 12, 79, 42, 68, 71, 77, 73]. At the core is the quantum dot, which contains a trapped electron(s) or hole(s) to form a physical qubit [28]. Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations. Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7], exploring different structures and material combinations.
Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and device availability. Major technological hurdles are related to the so-called interconnect bottleneck [68], and various fabrication challenges towards scaling up [10, 6, 19, 18]. On the upside, there have been significant efforts [38, 9, 68, 33, 23, 49, 52, 70] to tackle these challenges.
@@ III Problem Statement
The need to apply design space exploration (DSE) techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 65, 5]. These techniques or similar ones have been successfully applied across different levels of the quantum computing stack. For instance, they have been used to explore ion-trap quantum processors [45], evaluate compilation techniques [58] and qubit connectivities [41, 39, 20] for superconducting qubits. Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize. Starting from there [54, 43], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a Figure of Merit that combines different performance metrics, and iv) find models (behavioral, analytical, experimental data) that connect input parameters with metrics (or directly to the Figure of Merit). By using these DSE techniques, performance trends can be observed projecting current or future multidimensional design spaces, and points of optimality can be identified. From such an approach, it can be determined which architectural design choices (i.e., combinations of input parameter values) could provide better performance when benchmarked from small to large-scale quantum circuits.
@@
Although this methodology is well established and understood in many disciplines, it can be challenging to implement in such early stages for spin-qubit technologies. One of the most difficult aspects is the development of a simulation framework incorporating a range of representative architectural variables adopting a large-scale perspective. Firstly, it is difficult to predict which architectural features will be relevant for many generations of devices to come, and secondly, the framework itself has to be flexible to incorporate new ones easily.
In practice, this can be very time-consuming, and thus, typically, the design space is kept at a reasonable and, usually, small scope under specific technological assumptions. Considering the rapid progress of spin-qubits, the potential design space can be enlarged quickly with new considerations, rendering any previous exploration obsolete. Hypothesizing such a conventional DSE framework exists, it can be tedious to brute force thousands of designs and extensively analyze results. Hence, the relevancy of a framework automating such a DSE process, which can simplify performance evaluation of different spin-qubit architectures, is prominent.
@@
To capture the above problem statement, we ask the following four research questions:
@@
    Which optimization methods are best suited for automating the DSE per quantum cicruit? Answered in Section VII.1
@@
    Which architectural characteristics in the selected design space are key for building high-performance spin-qubit devices for specifically given quantum circuits? Answered in Section VII.2
@@ IV SpinQ and the design space
Table 1: Summary of values of the considered architecture variables.
To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables. These properties were selected to be representative of spin qubit devices as well as to create trade-offs between them. The latter will prevent maximizing one without causing a performance penalty through another (refer to Section V). The defined design space is summarized in Table 1, including their possible values. It should be noted that the DSE analysis in this work uses either discrete or categorical variables [46], as we aim to provide performance trends and rankings between optimization methods and architectures rather than absolute comparisons. An architecture is then characterized by one (valid) combination of variable values. Below, we describe the considered design space analytically.
@@ IV.1 Operational gate constraints
The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types can be executed simultaneously and (b) how many of them can be parallelized at the same time step. To highlight an example of a design trade-off, we note that the higher the degree of parallelization, the shorter the circuit depth will be, and therefore, the shorter the algorithm execution time, but it will result in higher crosstalk.
@@
    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X/Y gates can be executed simultaneously with Z gates.
@@
    These variables indicate how many gates, for each type, can be parallelized, expressed as a percentage relation to the total number of qubits. For example, xyD = 25 in a 40-qubit architecture means that at most ten X or Y gates can exist in a single cycle. Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates. The zD and sD variables indicate the same for Z rotations and shuttle gates, respectively. The tqgD variable indicates the parallelization amount for two-qubit gates. As these always involve two qubits, they are expressed as a percentage of half of the total number of qubits. A value of $-1$ means that the parallelization degree is not user-defined but dictated by a specific single-qubit gate implementation explained next.
@@
    This variable indicates in what manner single-qubit gates are carried out. A Global single-qubit gate implementation means the same rotation axis and angle are applied to all qubits. For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation. Then, the Semi-Global implementation is considered with the same rotation scheme as in [30] and [51], which is essentially equivalent to the implementations of the crossbar architecture proposed of [38]. In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [51]. A Local implementation allows for any parallelized combination of gates at any angle to be executed in a cycle. Finally, a Sequential single-qubit gate implementation does not allow any gate parallelization, meaning that every cycle will be occupied by one gate only. It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible. For instance, in the Global and Semi-Global implementations, the xyD variable is not applicable as these implementations predetermine the upper limit of single-qubit gates allowed during a cycle.
@@ IV.2 Implementation of Z rotations:
This boolean variable also stipulates the implementation of Z rotation gates on a hardware level. The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [51, 30, 44, 38]. This single-qubit manipulation belongs to the so-called “hopping spins” type [72], and it is possible to achieve an arbitrary rotation axis, but in this work, we only assume the aforementioned variety for simplicity. The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by magnetic pulse interactions. With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits. Similarly to the variables mentioned earlier, only certain combinations are possible depending on single_qubit_impl values. For example, with a Semi-Global single-qubit gate implementation and a pulse-based Z rotation gate, the zD variable is not applicable for the same reason the xyD variable is not applicable, as we explained before in category (b).
@@ IV.3 Connectivity of the device:
This variable stipulates the considered coupling graphs by addressing them through their average node degree. A coupling graph is a graph of possible direct interactions (edges) between qubits (nodes). The first coupling graph we used is a square grid (degree = 4). The second is identical to the first but adds one diagonal connection in each square (i.e., each qubit is coupled with six neighboring qubits and with three at the corners, degree = 6). Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8). Figure 1, presents the coupling graphs considered with color-coded edges for each of the three types. We have now reviewed all the variables related purely to hardware. Next, we will consider two variables associated also with the compilation process. More specifically, they change the routing process but based on the hardware communication constraints.
Figure 1: The different coupling graphs considered in the design space. The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8.
IV.4 Routing methods:

### v3

@@ Abstract
In the fast-paced field of quantum computing, identifying the architectural characteristics that will enable quantum processors to achieve high performance across a diverse range of quantum algorithms continues to pose a significant challenge. Given the extensive and costly nature of experimentally testing different designs, this paper introduces the first Design Space Exploration (DSE) for quantum-dot spin-qubit architectures. Utilizing the upgraded SpinQ compilation framework, this study explores a substantial design space comprising 29,312 spin-qubit-based architectures and applies an innovative optimization tool, ArtA (Artificial Architect), to speed up the design space traversal. ArtA can leverage 17 optimization configurations, significantly reducing exploration times by up to 99.1% compared to a traditional brute-force approach while maintaining the same result quality. After a comprehensive evaluation of best-matching optimization configurations per quantum circuit, ArtA suggests specific as well as universal architectural features that provide optimal performance across the examined circuits. Our work demonstrates that combining DSE methodologies with optimization algorithms can be effectively used to generate meaningful design insights for quantum processor development.
@@ I Introduction
As the field of quantum computing rapidly advances, different qubit technologies exhibit unique hardware and performance characteristics. It still remains to be seen which one (e.g., superconducting, trapped ions, quantum dots, photonics, defect-based on nitrogen-vacancy diamond centers) will succeed in scaling up quantum computing systems with high-quality qubits [66, 11]. Among them, spin qubits in quantum dots emerge as a compelling avenue for achieving scalability for practical quantum computation [11]. To this day, spin qubits are at an early stage in their development, with Intel’s Tunnel Falls chip currently boasting the highest count of twelve spin qubits [34]. Despite this, their inherent scalability advantages suggest that a robust two-dimensional design [78, 41, 32, 7, 91, 38, 77] could be scaled up relatively easily once fabrication techniques and quality improve up to a certain level. Designing a chip, however, involves multiple architectural design choices whose impact in the future can only be fully assessed through experimental studies post-scaling the technology. Exploring all these design possibilities simultaneously can be expensive, time-intensive, and impractical. Therefore, simulating how different design choices can affect performance in a range of quantum applications is crucial to assist the development. This process facilitates highlighting architectural insights that can guide the technology forward and will allow quantum researchers to make more informed decisions, streamline efforts, and hasten the development of these devices.
@@
Standard DSE processes [60, 47], however, will be impractical time-wise, especially for such large spaces, as they rely on a brute-force approach to traverse and subsequently analyze the design space. To address this, our approach employs ArtA (Artificial Architect), a built-in tool containing multiple optimization methods for automating the DSE process of SpinQ, thus taking less time than the brute-force approach with the same quality of results. The abilities of ArtA are (a) to suggest which of the seventeen optimization configurations can find the architecture with the desired ESP the fastest compared to the brute-force approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per quantum circuit.
Our results demonstrate ArtA’s ability to compare all optimization techniques for each quantum circuit and obtain a solution up to 99.1% faster, on average, than brute-forcing. Then, equipped with these insights, we move on to conduct a DSE analysis of $29,312$ spin qubit architectures and provide valuable insights into best design practices. Firstly, we highlight the critical need to maximize the parallelization of quantum gates rather than minimizing crosstalk between qubits. Secondly, we find that the communication method via shuttle operations is achieving higher ESP in large-scale circuits than SWAPs, but for single-qubit gates, pulse-based rotations are preferred over shuttle-based ones. Finally, a universal recommendation based on all tested circuits indicates that the likelihood of success increases when prioritizing the parallelization of single-qubit gates over two-qubit gates. This demonstrates that the combination of DSE techniques and optimization algorithms can effectively guide quantum processor designers with meaningful recommendations in the current and future stages of spin-qubit technology. Notably, we also observe that certain quantum circuits achieve optimal performance even with reduced hardware capabilities. This highlights the capability of our approach to uncover optimal architectural configurations without relying on highly complex or resource-intensive hardware designs.
The main contributions of this paper are:
@@
    The upgraded SpinQ, the first compiler and DSE framework for spin qubit architectures. In this version, we have updated the ESP formula to include operational, crosstalk, and decoherence-induced errors.
@@
The remainder of this paper is structured as follows: In Section II, we discuss the scalability potential of quantum-dot spin qubits as well as the importance of addressing engineering challenges while taking into account the performance of the resulting architecture. Then, in Section III, we motivate the need for an automated DSE analysis for spin qubit devices and formulate four research questions for this work. After that, in Section IV, we define the design space consisting of various input architectural variables and establish the new functionalities of SpinQ. Another upgrade in SpinQ is the enhanced ESP metric, described in Section V, which is used as the figure of merit for our exploration. In Section VI, we introduce ArtA and two new metrics to evaluate its performance. In the results Section VII, we first evaluate ArtA’s performance across all used quantum circuits and determine the best optimization method configuration. Then, we conduct a detailed analysis of the best architectural designs for each quantum circuit and form valuable insights. We conclude our work and provide ideas for future directions in Section VIII.
@@ II Challenges for scaling up spin qubit devices
Spin-qubit technologies are distinguished by their unique physical features, which position them as a highly scalable solution for quantum computing. The advantages of spin-qubits include a significantly smaller size — up to a thousand times less than other qubit technologies — combined with decades of semiconductor manufacturing expertise, long coherence times coupled with short gate durations, and high operational temperatures [87, 10, 32, 11, 92, 45, 78, 81, 88, 84]. At the core of this technology lies the quantum dot, which can contain a trapped electron(s) or hole(s) to form a physical qubit [28]. Spin qubits are manipulated electromagnetically using multiple precision-engineered gate electrodes that facilitate either single- or two-qubit operations through exact timing of pulse sequences across various quantum dot configurations. Studies have expanded these systems into one-dimensional and two-dimensional arrays [32, 7, 91, 38, 77], exploring different structures and material combinations.
Despite the mentioned advantages, spin qubit quantum processors are not as advanced as other qubit technologies in terms of qubit counts and device availability. Major technological hurdles are related to the so-called interconnect bottleneck [78], and various fabrication challenges towards scaling up [9, 6, 19, 18]. On the upside, there have been significant efforts [41, 8, 78, 33, 23, 54, 58, 80, 35] to tackle these challenges.
@@ III Problem statement
The need to apply DSE techniques for designing and optimizing full-stack quantum computing systems, or components thereof, has been emphasized in previous work [2, 74, 5]. These techniques and similar ones have been successfully applied across different levels of the quantum computing stack which predominantly rely on brute-force exploration (i.e., exhaustive search). This is largely due to the inherently small design spaces in current quantum computing systems, which renders brute-force evaluation computationally feasible. For instance, DSE methodologies have been used to explore ion-trap quantum processors [49], evaluate compilation techniques [64] and qubit connectivities [44, 42, 20] for superconducting qubits. Such techniques can abstractly convert architectural characteristics into design variables and, in this way, quantize the design space, expressing current and future design possibilities that otherwise would be impossible or too time-consuming to physically realize. Starting from there [60, 47], the steps for a proper DSE: i) Describe the problem in terms of input variables or parameters (design choices), ii) select the performance metrics, iii) choose a global cost function as a figure of merit that combines different performance metrics, and iv) find models (behavioral, analytical, from experimental data) that connect input parameters with metrics or directly to the figure of merit. Applying these DSE techniques allows researchers to uncover performance trends, model current and future multidimensional design spaces, and identify optimal design points across a wide range of application use cases.
@@
Although this methodology is well established and understood in many disciplines, it can be challenging to implement in such early stages for spin-qubit technologies. One of the most difficult aspects is the development of a simulation framework incorporating a range of representative architectural variables with a large-scale perspective. Firstly, it is difficult to predict which architectural features will be relevant for many generations of devices to come, and secondly, the framework itself has to be flexible to incorporate new ones easily.
In practice, performing exhaustive exploration can be highly time-consuming; thus, even in classical systems, the design space is typically constrained to a manageable size by applying simplifying technological assumptions. While optimization algorithms have been successfully employed to efficiently explore significantly larger design spaces in classical computing domains [73, 12, 46, 53, 90, 65, 68, 83], such techniques remain largely unexplored in quantum computing DSE applications. As spin-qubit technology rapidly advances, the design space continues to grow with new considerations, making earlier explorations quickly outdated. Even assuming a brute-force approach already exists, the exponential growth in possible configurations can soon make it impractical to evaluate thousands of designs. Therefore, there is a clear need for an automated framework capable of efficiently managing this DSE process, significantly reducing the exploration time across current and future spin-qubit architectures.
@@
To capture the challenges of the above problem statement, we ask the following four research questions:
@@
    Which optimization methods are most effective for automating the DSE process for each quantum circuit individually, as well as across all used circuits collectively? Answered in Section VII.1
@@
    Which architectural characteristics in the selected design space are key for building high-performance spin-qubit devices for the specifically given quantum circuits? Answered in Section VII.2
@@ IV SpinQ and the design space
Table 1: Summary of values of the considered architecture variables
To answer the first two research questions, we conducted an analysis of relevant spin-qubit architectural properties with a long-term view and distilled them into numerous variables. These properties were selected to be representative of spin qubit devices as well as to be able to form performance-related trade-offs between them. The latter will prevent “maximizing” one without causing a performance penalty through another (refer to Section V). The defined design space is summarized in Table 1, including their possible values. An architecture is then characterized by one (valid) combination of variable values. Below, we describe the considered design space analytically. It should be noted that the DSE analysis in this work uses either discrete or categorical variables [50], as we aim to provide performance trends and rankings between optimization methods and architectures.
@@ IV.1 Operational gate constraints
The set of gate constraints can be vast; therefore, we decided to abstract them into the following variables that affect which (a) gate types can be executed simultaneously and (b) how many of them can be parallelized at the same time step.
@@
    As the names suggest, these variables are boolean and signify whether specific gate types can be executed in parallel. For example, xy_z defines if X or Y gates can be executed simultaneously with Z gates.
@@
    These variables indicate how many gates of each type can be parallelized, expressed as a percentage relation to the total number of qubits. For example, xyD = 25 in a 40-qubit architecture means that at most ten X or Y gates can exist in a single cycle. Note that the notion of a cycle refers to the basic unit of time representing one step in a sequence of gates of a quantum circuit, and each step may contain multiple gates. The zD and sD variables indicate the same for Z rotations and shuttle operations, respectively. The tqgD variable indicates the parallelization amount for two-qubit gates. As these always involve two qubits, they are expressed as a percentage of half of the total number of qubits. A value of $-1$ means that the parallelization degree is not user-defined but dictated by a specific single-qubit gate implementation explained next.
@@
    This variable indicates in what manner single-qubit gates are carried out. A Global single-qubit gate implementation means the same rotation axis and angle are applied to all qubits. For qubits that do not participate in a cycle, locally applied disable instructions prevent their rotation. Then, the Semi-Global implementation is considered with the same rotation scheme as in [30] and [56], which is essentially equivalent to the implementations of the crossbar architecture proposed in [41]. In this implementation, qubit rotations are implemented semi-globally, meaning that either all qubits in odd or even column parities can be rotated at a time with the same axis and angle, and unwanted qubit rotations have to be reversed afterward by additional instructions [56]. A Local implementation enables arbitrary parallel execution of single-qubit gates, allowing any combination of rotation axes and angles to be applied simultaneously within a single cycle. Finally, a Sequential single-qubit gate implementation does not allow parallelizations, meaning that every cycle will be occupied by one gate only. It should be noted that depending on the single_qubit_impl value, only certain combinations of categories (a) and (b) are possible. For instance, in the Global and Semi-Global implementations, the xyD variable is not applicable as these implementations predetermine the upper limit of single-qubit gates allowed during a cycle.
@@ IV.2 Implementation of Z rotations:
This boolean variable stipulates the hardware-level implementation of Z rotation gates. The first option is a high-fidelity shuttle-based Z rotation, which is achieved with two time-sensitive qubit shuttles to and from a neighboring column [56, 30, 48, 41]. This single-qubit manipulation belongs to the so-called “hopping spins” type [82], and it is possible to achieve an arbitrary rotation axis, but in this work, we only assume the aforementioned variety for simplicity. The second option is a regular pulse-based Z rotation, where Z gates are implemented similarly to X/Y gates by magnetic pulse interactions. With the first choice being a unique characteristic of spin-qubits, we are interested in exploring which one of the two provides better performance across many quantum circuits. Similarly to the variables mentioned earlier, only certain combinations are possible depending on single_qubit_impl values. For example, with a Semi-Global single-qubit gate implementation and a pulse-based Z rotation gate, the zD variable is not applicable for the same reason the xyD variable is not applicable, as we explained before in category (b).
@@ IV.3 Connectivity of the device:
This variable stipulates the considered coupling graphs by addressing them through their average node degree. A coupling graph is a graph of possible direct interactions (edges) between qubits (nodes). The first coupling graph we used is a square grid (degree = 4). The second is identical to the first but adds one diagonal connection in each square (i.e., each qubit is coupled with six neighboring qubits and with three at the corners, degree = 6). Finally, the third adds the second diagonal in each square (i.e., each qubit is coupled with eight neighboring qubits and with three at the corners, degree = 8). Figure 1(a) presents the coupling graphs considered with color-coded edges for each of the three types. The actual dimensions of each coupling graph, and thus the overall size of the quantum architecture, are determined by the number of qubits required by the quantum algorithm under execution.
Figure 1: (a) The different coupling graphs considered in the design space. The coupling graph consists of the black line edges when degree = 4, black and green edges when degree = 6, and all edges when degree = 8 (b) Schematic overview of the crossbar architecture, taken from [56], with the various shared operational control lines: vertical (column line, CL), horizontal (row line, RL), and diagonal (qubit line, QL). These lines are used with precise pulse sequences and shared among multiple sites, sixteen quantum dots in this figure, to perform operations on qubits [41, 30, 48] under specific operational constants. Here, the qubits (green circles with numbers) are initialized in a checkerboard pattern. (c) Abstraction of the crossbar architecture taken from [55] representing the coupling graph between qubits with degree = 4. Each circle represents a quantum dot, and each edge represents a coupling link signifying allowed interactions.
To further ease the understanding of the architectural landscape of the design space, in Figure 1(b), we depict the crossbar architecture [41] as a real-world example of an architecture represented in the design space. In this figure, the operational lines confine sixteen quantum dots, eight of which are occupied by spin qubits. The operational gate constraints in this architecture arise from the shared-control lines and other physical properties as described in [41, 30, 48, 56]. Among these constraints, the coupling connectivity is particularly relevant, as it corresponds to a square grid with degree = 4 in this implementation. Consequently, the topology of the crossbar architecture can be abstracted as a two-dimensional grid, as depicted in Figure 1(c).
We have now reviewed all variables that pertain solely to the hardware. Next, we examine two additional variables that influence the compilation process. Specifically, these variables can alter the routing methodology based on the underlying hardware communication constraints.
IV.4 Routing methods via shuttling: